OpenStudy (anonymous):

Which is NOT a possible solution of the equation y = x – 3? (1, –2) (–4, –1) (3, 0) (6, 3)

OpenStudy (lgbasallote):

substitute x and y with the ordered pairs i.e. (1,-2) so x = 1 and y = -2 y = x - 3 substitute... -2 = 1 - 3 -2 = -2 therefore (1,-2) is a solution does that help?

OpenStudy (radar):

Semantics.......Since they are asking a NOT solution, Although that was a solution to the equations it is not a solution to the problem. lol